OpenStudy (anonymous):

Convert 200 cubic meters per second to cubic inches per minute.

OpenStudy (anonymous):

hmmm, i dont know....

OpenStudy (anonymous):

One meter = 3.2808 feet One m^3 = 3.2808^3 feet^3 one m^3 = 35.3134 feet ^3 200 x 35.3134 feet^3/m^3 = 7062.68 feet^3/sec 7062.68 x 12^3 = 12,204,303.5 in^3/sec 12,204,303.5 in^3/sec * 60sec/min = 732,258,210 in^3/min